OpenStudy (anonymous):

Ok, I need help on how to graph with Point Slope form. Like how to get something to Point slope form and then how to graph it.

OpenStudy (2bornot2b):

Suppose you are provided with a point, e.g. \((x_1,y_1)\), and a slope \(m\) so the line passing through that point and having that slope is \[y-y_1=m(x-x_1)\]

OpenStudy (2bornot2b):

What else do you want to know?

OpenStudy (anonymous):

Yes I know that but what about how to solve it and graph it

OpenStudy (2bornot2b):

Give me a problem..

OpenStudy (anonymous):

I'm not sure how, but I have this one I don't understand. Wrate an equation of the line that passes through (2,1) and is Parallel to the line Y=2x+7 A. y-2=2(x+1) B. y-1=1/2(x-2) C. y+1=1/2(x+2) D. y-1=2(x-1) How do I solve this?

OpenStudy (2bornot2b):

So your line passes through the point (2,1) and it is parallel to the line y=2x+7 which means your line should have a slope of 2 (when a line is expressed as y=mx+c, m is the slope of the line, here y=2x+7 therefore has a slope of 2, and since you want to find a line parallel to it, your line will also have a slope 2) So now the equation of your line would be \[y-1=2(x-2)\]

OpenStudy (anonymous):

Oh! Ok I kinda get it, but is the last 2 the 2 from the (2,1)

OpenStudy (2bornot2b):

You typed the option D wrong. It should be D. y-1=2(x-2)

OpenStudy (2bornot2b):

YES

OpenStudy (anonymous):

Ok! Thanks that makes so much sense to me now. But which letter would it be?

OpenStudy (2bornot2b):

It should be D, though you typed it wrong here. Check it, I think it should be y-1=2(x-2) instead of y-1=2(x-1)

OpenStudy (anonymous):

Its y-1=2(x-2)

OpenStudy (2bornot2b):

Then option D is correct.